# USC: Uncompromising Spatial Constraints for Safety-Oriented 3D Object Detectors in Autonomous Driving

### v4

@@ Abstract
We consider the safety-oriented performance of 3D object detectors in autonomous driving contexts. Specifically, despite impressive results shown by the mass literature, developers often find it hard to ensure the safe deployment of these learning-based perception models. Attributing the challenge to the lack of safety-oriented metrics, we hereby present uncompromising spatial constraints (USC), which characterize a simple yet important localization requirement demanding the predictions to fully cover the objects when seen from the autonomous vehicle. The constraints, as we formulate using the perspective and bird’s-eye views, can be naturally reflected by quantitative measures, such that having an object detector with a higher score implies a lower risk of collision. Finally, beyond model evaluation, we incorporate the quantitative measures into common loss functions to enable safety-oriented fine-tuning for existing models. With experiments using the nuScenes dataset and a closed-loop simulation, our work demonstrates such considerations of safety notions at the perception level not only improve model performances beyond accuracy but also allow for a more direct linkage to actual system safety.
@@ I Introduction
3D object detection is an essential function for scene understanding in autonomous driving (AD). It allows an autonomous vehicle (AV) to locate objects in a scene and classify them into different categories so that the AV can navigate through them effectively.
In the past years, computer vision and deep learning literature have contributed greatly to the growing performances of 3D object detectors. More recently, aiming at wider and safer deployment of AVs, several works have adapted the general literature and designed performance indicators that consider driving dynamics. For example, instead of the generic Intersection-over-Union (IoU) measure, the nuScenes benchmark uses translation errors (TE) to match the predictions and ground truths and calculates several true-positive metrics for more fine-grained evaluation [1]. While various driving-oriented metrics are proposed, one challenge still remaining is the definition of a straightforward criterion, i.e., what score a model should attain to be satisfactory and deployable. Lately, the mitigation of the challenge has become more urgent, as regulations such as the EU AI Act [2] require developers to demonstrate “an appropriate level of performance” before marketing their solutions.
@@
Hence, in this work, we consider tolerable errors for 3D object detectors from a safety point of view and propose uncompromising spatial constraints (USC) as potential bottom lines. Concretely, pivoting on object response and collision avoidance, we underpin a requirement that states an object detector’s prediction should fully cover its corresponding ground truth when seen from the AV’s location. The reason is that such full coverage shall (already) result in a low risk of collision at the object’s edges close to the AV. Fig. 1 shows an example scene and contrasts USC with existing practices, which generally demand perfect alignment between the prediction and the ground truth.
@@
To demonstrate the effect of our proposals, we conduct three sets of experiments with state-of-the-art models hosted on the MMDetection3D platform. First, we evaluate several models in the nuScenes dataset and check for safety concerns via USC-based metrics. We then improve one of the models through the safety-oriented loss function. Finally, by running the object detectors with an automatic emergency braking (AEB) function in a closed-loop setup, we show a better correlation with the AV’s collision rates enabled by USC. Altogether, our work serves as an initiative to address safety principles and remedies for imperfect 3D object detectors in autonomous driving systems.
@@ II Related Work
Object detection has attracted great research attention over the past two decades. Readers may refer to a recent survey that delineates the development [5]. In the following, we focus on the studies that concern AD and safety.
@@
As one of the early results, Cheng et al. proposed a conceptual safety-aware hardening framework for object detectors [6]. Specifically, the authors defined a distance-based criticality criterion and adapted the NN architecture of a baseline model to enhance its performance in critical zones. Based on a similar idea, Lyssenko et al. refined the definition of criticality by considering object dynamics and improved pedestrian detection accordingly [7]. While these results contributed to better classification and detection of safety-critical objects generally, our study complements them by precising localization requirements at a finer scale. In fact, the introduced nuScenes benchmark lies in this direction [1]. It adopts the NuScenes Detection Score (NDS), computing not only the conventional mean average precision (mAP) metric [5] but also true-positive error measures on the localization function. Concurrently, Philion et al. proposed planner-centric metrics, which compute the KL-divergence between the motion planning outcomes (i.e., trajectory waypoints) based on predictions and ground truths respectively [8]. Although the latter approach appears plausible, a recent validation work interestingly found NDS better correlating with actual AV collision rates, indicating the advantage of true-positive error measures at the object detection model level [9]. Hence, our investigation follows this direction, but as we shall see, while NDS demands perfectly aligned predictions, our work focuses on safety-oriented principles.
@@
In the most recent literature, there exist several similar efforts. For instance, Deng et al. proposed Support Distance Error (SDE), computing the marginal gaps on the edges of a prediction and its targeted ground truth [10]. Their proposal, however, is tailored for a point cloud-based object representation in the BEV plane. Similarly, Mori et al. considered several safety-oriented adaptations for existing evaluation protocols, e.g., by redefining the shortest distance between two bounding boxes [11]. They also derived pass/fail criteria based on human perception performance. While the analysis with human benchmarks has been shown feasible, some considerations may be overly complicated, as reported in the work itself. Henceforth, we address localization constraints that are simplistic and general for 3D object detectors using the typical bounding box-based representation.
@@
Lastly, there have also been interesting results that compensate for the imperfection of learning-based models. For example, De Gerancey et al. leveraged conformal prediction to post-process the predicted bounding boxes and obtain a statistical performance guarantee in ground truth alignment [3]. Similarly, Schuster et al. analytically derived a minimum multiplicative factor for enlarging the predictions so as to properly enclose the ground truths [4]. While these results focus on the 2D image plane, our work distinctively targets 3D object detection, which is more directly linked with AD applications, and improves the models in a continual learning fashion.
III Preliminaries
As the foundation of our work, we further introduce the mostly adopted practices for object detector evaluation, including the typical IoU-based approach [5] and nuScenes’ protocol [1].
For 3D object detection in AD, the common practice is to abstract objects with upright 3D bounding boxes [5]. It also usually assumes the AV at the origin $O(0,0,0)$ and a right-handed coordinate system with the $z$-axis parallel to the AV’s heading and $x$ pointing to the right. Then, an object’s bounding box, taken as the ground truth, is annotated by a tuple $\hat{\mathbf{G}}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=\,}}}(x_{\mathbf{G}},y_{\mathbf{G}},z_{\mathbf{G}},l_{\mathbf{G}},h_{\mathbf{G}},w_{\mathbf{G}},\theta_{\mathbf{G}})$. Here, $x_{\mathbf{G}}$, $y_{\mathbf{G}}$, $z_{\mathbf{G}}$ are the center coordinates, $l_{\mathbf{G}}$ is the length (parallel with the $x$-axis when the orientation of the box is 00), $h_{\mathbf{G}}$ is the height (along the $y$-axis), $w_{\mathbf{G}}$ is the width (along the $z$-axis), and $\theta_{\mathbf{G}}$ is the rotation around the $y$-axis. Similarly, a prediction given by the object detector is represented by $\hat{\mathbf{P}}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=\,}}}(x_{\mathbf{P}},y_{\mathbf{P}},z_{\mathbf{P}},l_{\mathbf{P}},h_{\mathbf{P}},w_{\mathbf{P}},\theta_{\mathbf{P}})$.
Now, for assessing a prediction against a given ground truth, most studies use the IoU measure [5]. For its calculation, the 3D cuboids of the prediction and the ground truth, denoted as $\mathbf{P}\subset\mathbb{R}^{3}$ and $\mathbf{G}\subset\mathbb{R}^{3}$, are first reconstructed from the tuple representations through their corners. Then, IoU is defined as:
$\mathsf{IoU}(\mathbf{P},\mathbf{G})\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=\,}}}\frac{\mathsf{Vol}(\mathbf{P}\cap\mathbf{G})}{\mathsf{Vol}(\mathbf{P})+\mathsf{Vol}(\mathbf{G})-\mathsf{Vol}(\mathbf{P}\cap\mathbf{G})}\,,$ (1)
where $\mathsf{Vol}(\cdot)$ is the volume of a given polyhedron, and $\mathbf{P}\cap\mathbf{G}$ denotes the intersection of the prediction $\mathbf{P}$ and ground truth $\mathbf{G}$. For $\mathbf{P}$ and $\mathbf{G}$, the volume can be easily computed with their dimensions (i.e., length, width, and height). For the intersection $\mathbf{P}\cap\mathbf{G}$, one normally derives its projected area in the BEV plane as well as its height along the $z$-axis separately and then multiplies them together to get the volume. In practice, there are advanced algorithms and software packages such as PyTorch3D [12] to implement the functions.
With the IoU’s definition, more conventional protocols traverse the object classes and assess the set of predictions against all the ground truths per class [5]. Then, based on given IoU thresholds, the predictions are matched with the ground truths and grouped into true positives (TPs), false positives (FPs), and false negatives (FNs), whose counts are in turn used to compute recall, precision, and the mean Average Precision (mAP) metric. By contrast, the nuScenes benchmark adopts thresholds based on BEV center distances when calculating mAP and defines the more fine-grained NuScenes Detection Score (NDS) as follows:
$\mathsf{NDS}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=\,}}}\frac{1}{10}[5\cdot\mathsf{mAP}+\sum_{\mathsf{TP}\in\mathbb{TPS}}(1-\min(1,\mathsf{TP}))],$ (2)
where $\mathbb{TPS}$ denotes the set of five TP measures for matched predictions and ground truths, concerning translation, orientation, scale, velocity, and attribute predicting errors. To be specific, the BEV center distance and the TP translation error (TE) between a prediction and a ground truth can be calculated with their representation attributes:
$\mathsf{TE}(\hat{\mathbf{P}},\hat{\mathbf{G}})=\sqrt{(x_{\mathbf{P}}-x_{\mathbf{G}})^{2}+(z_{\mathbf{P}}-z_{\mathbf{G}})^{2}}.$ (3)
For each class, TE is computed for all matched pairs and summarized into an average TE (ATE) score. Other TP measures follow the same process. For their definitions, readers are deferred to the nuScense benchmark [1].
As introduced, the IoU measure provides a generic indication of how well the prediction is aligned with the ground truth. Likewise, while NDS undertakes several driving-related considerations, it still regards errors from different directions equally and requires perfect predictions. We now present our safety-oriented spatial constraints which are motivated by error-tolerating considerations for collision avoidance from the AV’s perspective.
IV Safety-Oriented Spatial Constraints
@@
Figure 2: We transform 3D bounding boxes into the PV and BEV planes for establishing USC and corresponding quantitative measures.
@@
To iterate, our goal is to formulate uncompromising spatial constraints (USC) that ensure a given prediction $\mathbf{P}$ fully covers the ground truth $\mathbf{G}$ from the AV’s perspective. Technically, as Fig. 2 depicts, we transform them into the PV and BEV planes, which are easier to handle than the direct 3D world. In the following Sec. IV-A, we first address PV which naturally reflects driving situations. Then, as PV foregoes the important depth information, we consider additional BEV constraints in Sec. IV-B. Finally, Sec. IV-C consolidates the respective PV and BEV constraints for comprehensive model evaluation.
IV-A Box Enclosure in the PV Plane
@@
To formulate the spatial constraint in the PV plane, we first have to take a perspective projection for the prediction $\mathbf{P}$ and the ground truth $\mathbf{G}$. This can be done straightforwardly for camera-based object detectors with the camera parameters and images. As for lidar-based models, one may borrow the camera system in the sensor suite or simply apply a pinhole-based perspective projection using a focal length $f=1$. This is because our following analysis only considers the relative position and size of the projected prediction and ground truth. As long as the same perspective projection (with any $f$) is applied to both the prediction $\mathbf{P}$ and the ground truth $\mathbf{G}$, we can examine their relation in the PV plane. To elaborate, for a point $p(x_{p},y_{p},z_{p})$ in the 3D world, one can obtain its coordinates $(a,b)$ in the PV plane with the projection formula [13], written as:
@@
As studied in [3, 4], we can now formulate a direct constraint in the PV plane that requires the prediction $\mathbf{P}^{\mathcal{PV}}$ to fully enclose the ground truth $\mathbf{G}^{\mathcal{PV}}$, i.e.,
Intuitively, without a proper enclosure, the AV risks a higher chance of bumping into the object on its sides. For a quantitative characterization, the previous work [3, 4] adopted the IoU function in 2D. However, we find Intersection-over-Ground-Truth (IoGT) more suitable here:
with $\mathsf{Area}(\cdot)$ computing the area of a given polygon. More concretely, Fig. 3 illustrates our rationale. It can be observed that while IoU indicates the degree of alignment, IoGT reflects the degree of enclosure and better serves the purpose of the PV constraint. Consequently, IoGT maximizes at $1$ when the prediction $\mathbf{P}^{\mathcal{PV}}$ fully encloses the ground truth $\mathbf{G}^{\mathcal{PV}}$.
Now, with the PV constraint and its quantitative measure formulated, there is another risk that may be neglected: Even when $\Pi^{\mathcal{PV}}$ in Eq. (5) is satisfied, it is still possible that in the 3D world, the ground truth $\mathbf{G}$ is not fully covered by the prediction $\mathbf{P}$ from the AV’s location. For instance, the prediction $\mathbf{P}$ might be farther from the AV yet much larger than the ground truth $\mathbf{G}$, so that after the perspective projection, $\mathbf{P}^{\mathcal{PV}}$ indeed encloses $\mathbf{G}^{\mathcal{PV}}$. To prevent such cases, we need to consider the depth information, or distance underestimation to be specific, in the BEV plane.
IV-B Distance Underestimation in the BEV Plane
As motivated, a rule of thumb in the BEV plane is that the prediction should be no farther than the ground truth. We provide a precise definition and a quantitative measure in the following.
First, as in the PV plane, we need the projected bounding boxes in the BEV plane. To such end, we perform an orthographic projection on the prediction $\mathbf{P}$ and the ground truth $\mathbf{G}$. Since we have assumed upright 3D bounding boxes, an orthographic projection simply boils down to taking the BEV attributes from the 3D representations, e.g., $(x_{\mathbf{G}},z_{\mathbf{G}},l_{\mathbf{G}},w_{\mathbf{G}},\theta_{\mathbf{G}})$ for the ground truth $\mathbf{G}$, and then the BEV bounding boxes can be reconstructed. In the following, we denote the projected prediction as $\mathbf{P}^{\mathcal{BEV}}$ and ground truth as $\mathbf{G}^{\mathcal{BEV}}$.
@@ IV-B Distance Underestimation in the BEV Plane
That said, we note a pitfall where $\Pi^{\mathcal{BEV}}_{c}$ is satisfied but some part of the ground truth $\mathbf{G}^{\mathcal{BEV}}$ is still exposed, as depicted in Fig. 4. It shows that for full coverage over the ground truth $\mathbf{G}^{\mathcal{BEV}}$, one has to consider the entire side(s) facing the AV. Hence, we utilize the extreme points of $\mathbf{P}^{\mathcal{BEV}}$ and $\mathbf{G}^{\mathcal{BEV}}$ from the AV’s perspective and further ensure there is no intersection in these sides. Formally, we define now the BEV constraint as:
@@
For the quantitative measure in the BEV plane, we observe generally that the farther the prediction is placed from the AV in comparison to the ground truth, the larger a portion of the ground truth is exposed. Therefore, we leverage the defined representative points (i.e., the closest, left-most, and right-most points from the AV’s perspective) and formulate an average distance ratio (ADR) as follows:
@@
Essentially, we compute individually how far the prediction’s representative points are from the AV compared to their ground truth counterparts and aggregate the point-wise ratio by taking the geometric mean. Here, using the geometric mean instead of other central tendencies, such as the arithmetic mean, gives more robust results, e.g., in the case of a large prediction [16]. Lastly, similar to IoGT, which maximizes when the prediction arrives at a “safe” state, ADR also saturates at $1$ if the prediction’s points are all closer than their corresponding ground truth points.
IV-C Consolidation for Model Evaluation

### v5

@@ Abstract
In this work, we consider the safety-oriented performance of 3D object detectors in autonomous driving contexts. Specifically, despite impressive results shown by the mass literature, developers often find it hard to ensure the safe deployment of these learning-based perception models. Attributing the challenge to the lack of safety-oriented metrics, we hereby present uncompromising spatial constraints (USC), which characterize a simple yet important localization requirement demanding the predictions to fully cover the objects when seen from the autonomous vehicle. The constraints, as we formulate using the perspective and bird’s-eye views, can be naturally reflected by quantitative measures, such that having an object detector with a higher score implies a lower risk of collision. Finally, beyond model evaluation, we incorporate the quantitative measures into common loss functions to enable safety-oriented fine-tuning for existing models. With experiments using the nuScenes dataset and a closed-loop simulation, our work demonstrates such considerations of safety notions at the perception level not only improve model performances beyond accuracy but also allow for a more direct linkage to actual system safety.
© 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
3D object detection is essential for scene understanding in autonomous driving (AD). It allows an autonomous vehicle (AV) to locate objects in a scene and classify them into different categories so that the AV can navigate through them effectively.
In the past years, computer vision and deep learning literature have contributed greatly to the growing performances of 3D object detectors. More recently, aiming at broader and safer deployment of AVs, several works have adapted the general literature and designed performance indicators that consider driving dynamics. For example, instead of the generic Intersection-over-Union (IoU) measure [1], the nuScenes benchmark uses translation errors (TE) to match the predictions and ground truths and calculates several true-positive metrics for more fine-grained evaluation [2]. While various driving-oriented metrics are proposed, a remaining challenge is the definition of a straightforward criterion, i.e., what score a model should attain to be satisfactory and deployable. Lately, the mitigation of the challenge has become more urgent, as regulations such as the EU AI Act [3] require developers to demonstrate “an appropriate level of performance” before marketing their solutions.
@@
Hence, in this work, we consider tolerable errors for 3D object detectors from a safety point of view and propose uncompromising spatial constraints (USC) as potential bottom lines. Concretely, pivoting on object response and collision avoidance, we underpin a requirement as follows: An object detector’s prediction should fully cover its corresponding ground truth when seen from the AV’s location. The reason is that such full coverage should already result in a low risk of collision at the object’s edges close to the AV. Fig. 1 shows an example scene and contrasts USC with existing practices, which generally demand perfect alignment between the prediction and the ground truth.
@@
To demonstrate the efficacy of our proposals, we conduct three sets of experiments with state-of-the-art 3D object detectors. First, we evaluate the models using the nuScenes dataset [2] and check for safety concerns via USC-based metrics [^1] The utilization of the nuScenes datasets in this paper is for knowledge dissemination and scientific publication and is not for commercial use.. We then improve one of the models through the safety-oriented loss function. Finally, by running the object detectors with an automatic emergency braking (AEB) function in closed-loop simulations, we show that USC enables a better correlation between component-level metrics and the AV’s collision rates. Altogether, our work serves as an initiative to address safety principles and remedies for imperfect 3D object detectors in autonomous driving systems.
@@ II Related Work
Object detection has attracted great research attention over the past two decades. Readers may refer to a recent survey that delineates the development [6]. In the following section, we focus on studies that are concerned with AD and safety.
@@
As one of the early results, Cheng et al. proposed a conceptual safety-aware hardening framework for object detectors [7]. Specifically, the authors defined a distance-based criticality criterion and adapted the NN architecture of a baseline model to enhance its performance in critical zones. Based on a similar idea, Lyssenko et al. refined the definition of criticality by considering object dynamics and improved pedestrian detection accordingly [8]. While these results contributed to better recognition of safety-critical objects in general, our study complements them by precising localization requirements at a finer scale. In fact, the introduced nuScenes benchmark lies in this direction [2]. It adopts the NuScenes Detection Score (NDS), computing not only the conventional mean average precision (mAP) metric [6] but also true-positive error measures on the localization function. Concurrently, Philion et al. proposed planner-centric metrics, which compute the KL-divergence between the motion planning outcomes (i.e., trajectory waypoints) based on predictions and ground truths respectively [9]. Although the latter approach appears plausible, a recent validation work interestingly found NDS better correlating with actual AV collision rates, indicating the advantage of true-positive error measures at the object detection model level [10]. Hence, our investigation follows this direction, but as we shall see, while NDS demands perfectly aligned predictions, our work focuses on safety-oriented principles.
@@
In the most recent literature, several similar efforts have been made. For instance, Deng et al. proposed Support Distance Error (SDE), computing the marginal gaps on the edges of a prediction and its targeted ground truth [11]. Their proposal, however, is tailored for a point cloud-based object representation in the BEV plane. Similarly, Mori et al. considered several safety-oriented adaptations for existing evaluation protocols, e.g., by redefining the shortest distance between two bounding boxes [12]. They also derived pass/fail criteria based on human perception performance. While the analysis with human benchmarks has been shown feasible, some assumptions may be overly complicated, as reported in the work itself. Henceforth, we address localization constraints that are simplistic and general for 3D object detectors using the typical bounding box-based representation.
@@
Lastly, there have also been interesting results that compensate for the imperfection of learning-based models. For example, De Gerancey et al. leveraged conformal prediction to post-process the predicted bounding boxes and obtain a statistical performance guarantee in ground truth alignment [4]. Similarly, Schuster et al. analytically derived a minimum multiplicative factor for enlarging the predictions so as to properly enclose the ground truths [5]. While these results focus on the 2D image plane, our work distinctively targets 3D object detection, which is more directly linked with AD applications and improves the models in a continual learning fashion.
III Safety-Oriented Spatial Constraints
@@
Figure 2: We transform 3D bounding boxes (BBs) onto the PV and BEV planes for establishing USC and corresponding quantitative measures. Gold: original BB; Blue: converted PV or BEV BB.
@@
To iterate, using the standard upright 3D bounding box representation [1, 2], our goal is to formulate uncompromising spatial constraints (USC) that ensure a given prediction $\mathbf{P}$ fully covers the ground truth $\mathbf{G}$ when seen by the AV. Technically, as Fig. 2 depicts, we transform them into the PV and BEV planes, which are easier to handle than the direct 3D world. In the following Sec. III-A, we first address PV, which naturally reflects driving situations. Then, as PV foregoes the important depth information, we consider additional BEV constraints in Sec. III-B. Finally, Sec. III-C consolidates both PV and BEV constraints for comprehensive model evaluation.
III-A Box Enclosure in the PV Plane
@@
To formulate the spatial constraint in the PV plane, we first have to take a perspective projection for the prediction $\mathbf{P}$ and the ground truth $\mathbf{G}$. This can be done straightforwardly for camera-based object detectors with the camera parameters and images. As for lidar-based models, one may borrow the camera system in the sensor suite or simply apply a pinhole-based perspective projection using a focal length $f=1$. This is because our following analysis only considers the relative position and size of the projected prediction and ground truth. As long as the same perspective projection (with any $f$) is applied to both the prediction $\mathbf{P}$ and the ground truth $\mathbf{G}$, we can examine their relation in the PV plane. To elaborate, for a point $p\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=\,}}}(x_{p},y_{p},z_{p})$ in the 3D world, one can obtain its coordinates $(a,b)$ in the PV plane with the projection formula [13], written as:
@@
Similar to the prior work [4, 5], we now formulate a direct constraint in the PV plane that requires the prediction $\mathbf{P}^{\mathcal{PV}}$ to fully enclose the ground truth $\mathbf{G}^{\mathcal{PV}}$, i.e.,
Intuitively, without a proper enclosure, the AV risks a higher chance of bumping into the object on its sides. For a quantitative characterization, however, the previous work [4, 5] adopted the IoU function in 2D. On the contrary, we find Intersection-over-Ground-Truth (IoGT) more suitable here:
with $\mathsf{Area}(\cdot)$ computing the area of a given polygon. Fig. 3 illustrates our rationale. Concretely, while IoU indicates the degree of alignment, IoGT reflects the degree of enclosure and better serves the purpose of the PV constraint. In addition, IoGT maximizes at $1$ when the prediction $\mathbf{P}^{\mathcal{PV}}$ fully encloses the ground truth $\mathbf{G}^{\mathcal{PV}}$.
Now, with the PV constraint and its quantitative measure formulated, there is another risk that may be neglected: Even when $\Pi^{\mathcal{PV}}$ in Eq. (2) is satisfied, it is still possible that in the 3D world, the ground truth $\mathbf{G}$ is not fully covered by the prediction $\mathbf{P}$ from the AV’s location. For instance, the prediction $\mathbf{P}$ might be farther from the AV yet much larger than the ground truth $\mathbf{G}$, so that after the perspective projection, $\mathbf{P}^{\mathcal{PV}}$ indeed encloses $\mathbf{G}^{\mathcal{PV}}$. To prevent such cases, we need to consider the depth information or, more specifically, distance underestimation in the BEV plane.
@@
III-B Distance Underestimation in the BEV Plane
As motivated, a rule of thumb in the BEV plane is that the prediction should be no farther than the ground truth. We formulate such a rule and its quantitative measure in the following.
First, as in the PV plane, we need the projected bounding boxes in the BEV plane. To such end, we perform an orthographic projection on the prediction $\mathbf{P}$ and the ground truth $\mathbf{G}$. Since we have assumed upright 3D bounding boxes, an orthographic projection simply boils down to taking the rectangle on the BEV plane, as Fig. 2 shows. In the following, we denote the projected prediction as $\mathbf{P}^{\mathcal{BEV}}$ and ground truth as $\mathbf{G}^{\mathcal{BEV}}$.
@@ III-B Distance Underestimation in the BEV Plane
That said, as depicted in Fig. 4, we note a pitfall where $\Pi^{\mathcal{BEV}}_{c}$ is satisfied but some part of the ground truth $\mathbf{G}^{\mathcal{BEV}}$ is still exposed. For full coverage over the ground truth $\mathbf{G}^{\mathcal{BEV}}$, one has to consider the entire side(s) facing the AV. Hence, we utilize the extreme points of $\mathbf{P}^{\mathcal{BEV}}$ and $\mathbf{G}^{\mathcal{BEV}}$ from the AV’s perspective and further ensure there is no intersection in these sides. Formally, we now define the BEV constraint as:
@@
For the quantitative measure in the BEV plane, we observe that, in general, the farther the prediction is placed from the AV in comparison to the ground truth, the larger a portion of the ground truth is exposed. Therefore, we leverage the defined representative points (i.e., the closest, left-most, and right-most points from the AV’s perspective) and formulate an average distance ratio (ADR) as follows:
@@
Essentially, we compute individually how far the prediction’s representative points are from the AV compared to their ground truth counterparts and aggregate the point-wise ratio by taking the geometric mean. Here, using the geometric mean instead of other central tendencies, such as the arithmetic mean, gives more robust results, e.g., in the case of a large prediction [16]. Lastly, similar to IoGT, which maximizes when the prediction arrives at a “safe” state, ADR also saturates at $1$ if the prediction’s representative points are all closer than their corresponding ground truth points.
III-C Consolidation for Model Evaluation